### v1

# IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTION

Abstract
Generative models have shown significant achievements in audio generation tasks. However, existing models struggle with complex and detailed prompts, leading to potential performance degradation. We hypothesize that this problem stems from the low quality and relatively small quantity of training data. In this work, we aim to create a large-scale audio dataset with rich captions for improving audio generation models. We develop an automated pipeline to generate detailed captions for audio-visual datasets by transforming predicted visual captions, audio captions, and tagging labels into comprehensive descriptions using a Large Language Model (LLM). We introduce Sound-VECaps, a dataset comprising M high-quality audio-caption pairs with enriched details including audio event orders, occurred places and environment information. We demonstrate that training with Sound-VECaps significantly enhances the capability of text-to-audio generation models to comprehend and generate audio from complex input prompts, improving overall system performance. Furthermore, we conduct ablation studies of Sound-VECaps across several audio-language tasks, suggesting its potential in advancing audio-text representation learning. Our dataset and models are available at https://yyua8222.github.io/Sound-VECaps-demo/.
Index Terms— Audio generation, audio retrieval, diffusion model, audio-language dataset
1 Introduction
Recent advances in generative modelling have prompted text-to-audio generation models to achieve performance improvements. In particular, the development of fundamental language model [1, 2] and sampling-based diffusion model [3, 4] have driven the creation of powerful and robust systems [5, 6, 7] that can generate correlated and high-fidelity audio clips.
Despite achievements in generating audio with simple and straightforward captions, current models struggle with more complex prompts that present detailed information, which is commonly referred to the challenge as “prompt following ” [8]. We hypothesize that the fundamental reason for this limitation is the simplicity of current training datasets. As highlighted in Table 1, existing audio-caption datasets often fall short in both quality and detailed information. The captioning procedures of these datasets mainly rely solely on audio, offering simple captions that average fewer than words per sentence. This audio-only approach restricts the ability to provide more useful and enriched content. Therefore, the lack of information beyond basic prompts or sound tags restricts the system’s ability to improve its performance on complex and detailed prompts. In addition, the simplicity of the caption often results in situations where a single prompt corresponds to multiple audio files, causing instability in the quality and relevance of the generated outputs. In contrast, visual features have been proved to convey more information. Although the latest large-scale dataset, Auto-ACD [9], applied video features through automatic captioning pipelines to enhance the caption (18 words per caption), it only accounted for a single frame feature, and the prompt has been designed to avoid the visual information.
| Dataset | Number | Avg. Len | Loc. Inf | Env. Inf |
| AudioSet [10] | M | Label | Label | |
| Clotho [11] | K | K | K | |
| AudioCaps [12] | K | K | K | |
| WavCaps [13] | K | K | K | |
| Auto-ACD [9] | M | M | K | |
| Sound-VECaps | M | M | M | |
| Sound-VECaps | M | M | M |
In this paper, we aim to leverage external visual information to enhance the captions of the training dataset, guiding the system to follow the prompts. The integration of visual and audio features in the caption is designed to train the system to mimic human-like behaviour by constructing complete scenarios within a scene, thereby enhancing the alignment between the prompt and the sound. Specifically, we propose a new pipeline to construct a large-scale audio-language dataset with vision-enhanced captions. Our approach first involves collecting external visual information using state-of-the-art (SoTA) image captioning models. These visual captions, combined with audio information, are then used to create new, enriched captions through Large Language Models (LLMs). By incorporating additional visual information, our method ensures the accuracy of audio details while enhancing the captions with comprehensive content, including temporal, spatial, and contextual elements related to the environment. Building on the existing largest audio-visual dataset, AudioSet [10], we introduce a new dataset comprising over million audio-caption pairs, named Sound-VECaps. This dataset leverages the enhanced capabilities of vision-language integration to produce more detailed and contextually rich audio captions.
Using Sound-VECaps as the training dataset, our experiments with the audio generation model, AudioLDM [14], show substantial improvements over baseline models. In addition, we propose a new benchmark for text-to-audio generation by constructing an enhanced AudioCaps [12] testing set with human-involved supervision (same audio with better prompts). Specifically, the AudioLDM-L model trained on Sound-VECaps achieves a Frechet Audio Distance (FAD) score of on the AudioCaps dataset. It further improves to a score of on the AudioCaps-Enhanced testing set, significantly outperforming current SoTA models. Moreover, we conduct experiments on Sound-VECaps across various audio-language tasks, demonstrating that systems trained on Sound-VECaps also achieve SoTA performance in several specific audio-domain tasks, such as audio retrieval. Nevertheless, we investigate the effectiveness of the visual-only content within the caption and the influence of these features during inference. An external version of Sound-VECaps that excludes all the visual-only information (Sound-VECaps) is also provided for different research purposes.
The remainder of this paper is organized as follows. Section 2 introduces the generation pipeline of the proposed dataset, followed by the introduction of the audio generation system which is trained on our datasets in Section 3. Section 4 presents the experimental results which evaluate the re-captioned dataset on audio generation as well as several audio-language tasks. Conclusions are given in Section 5.
2 Dataset
As the generation pipeline displayed in Figure 1, the prompt that LLMs used to construct the caption consists of three different information, visual captions from the video, audio captions from the waveform, and the label taggings provided by the original dataset.
2.1 Visual information Captioning
We assume that video features can provide more guidance on the events, such as the position, ordering, acoustic environment and some other details of the audio events. Different from other captioning approaches [9] that involve a single frame of the whole video, we utilize the complete video information to enable the model to gain a more detailed understanding of each occurring event. However, SoTA audio captioning systems [15, 16] follow a similar strategy to understand the whole video clip by averaging the visual feature of the frames into a single feature dimension, which loses the temporal information. Hence, we apply image captioning models to capture the visual information for each frame to keep the temporal information. For instance, we follow the visual-caption generation pipeline of Stable Diffusion 3 [17] and apply the CogVLM [18] as the visual feature extraction model. To improve the efficiency, we clip each video by second and generate the caption of each frame through CogVLM, e.g., captions for eleven frames for a 10-second audio.
2.2 Audio information Captioning
Unlike other modalities, the audio feature is sometimes not highly correlated to visual information, such as background sound sources and invisible sound sources. Two constraints are provided as the audio feature for re-captioning to guide the LLM to generate the captions related to the actual auditory information. One is the label feature provided by the original AudioSet dataset, another is a simple audio caption generated by the audio captioning model. In our system, we applied the SoTA captioning model, EnCLAP [19], to generate concise and brief captions.
2.3 Caption Generation
2.4 Dataset Processing
Due to the issues of some videos being too old or concerns related to privacy, we collected a total of million videos from the AudioSet dataset. In addition, around k video clips are skipped during the captioning pipeline due to the sensitive policy of the LLMs (e.g., violence). Furthermore, we found that some video clips present static visual information and the auditory features are usually background sounds. Hence, a filtering strategy is developed to avoid the captions for static video scenes (more than % same frames in each video) and improve the correctness of the dataset. Overall, we construct the proposed Sound-VECaps consisting of M audio-caption pairs. The Sound-VECaps provides two different versions of captions for different purposes, specifically, Sound-VECaps containing only audible contents or environmental-descriptive information, and Sound-VECaps with full detailed information including visual features, e.g., texts, names, and colours.
| Model | Training Dataset | AudioCaps | AudioCaps-Enhanced | Best Result | |||||
| KL | IS | FAD | KL | IS | FAD | (%) | MOS | ||
| AudioGen [21] | AC+AS+8 others | 9.93 | |||||||
| AudioLDM [14] | AC+AS+2 others | ||||||||
| Tango2 [6] | AudioCaps | ||||||||
| AudioLDM2-Large [5] | AC+AS+6 others | 1.22 | |||||||
| AudioLDM-Proposed | Sound-VECaps | ||||||||
| AudioLDM-Proposed-L | Sound-VECaps | 1.49 | 1.17 | 7.96 | 1.06 | 43.59 | 4.05 | ||
3 System Architecture
To evaluate the impact of the enhanced prompt, we conduct experiments on basic AudioLDM [14] models. AudioLDM uses an encoder-decoder architecture to compress the waveform and a diffusion-based generative module to produce the audio features. For instance, the system can be divided into four sections: a CLAP encoder for condition embedding, a latent diffusion-based model to generate audio features within the latent space, a variational autoencoder (VAE) decoder to reconstruct the information into a mel spectrogram, and a generative adversarial network (GAN) vocoder [22] to produce the waveform as the final output.
Instead of using CLAP [1] for audio and text feature embedding during the training and inference stages respectively, our experiment only takes caption as the condition throughout the whole stage. Thus the CLAP encoder is replaced with T5 [2] text encoder to understand the textual information better. Following the idea of taking the T5 embedding in various lengths with masking [23], a cross-attention module is applied after each convolution layer of the UNet structure to undertake the condition. The rest modules exactly follow the design of AudioLDM and our system takes the pre-trained VAE decoder and Hifi-GAN vocoder for audio feature reconstruction.
4 Experiments
4.1 Evaluation Dataset
Despite the evaluation procedure on AudioCaps testing set through previous generative models, we introduce a novel benchmark for audio generation with enriched and enhanced captions. Utilizing the same audio samples, we apply the proposed re-captioning pipeline to generate improved captions for the proposed testing dataset. Specifically, human supervision is integrated into the captioning process to ensure the accuracy and semantic relevance of each prompt. Following the structure of the original AudioCaps testing set, the proposed AudioCaps-Enhanced testing dataset includes five different captions for each audio clip, totalling captions for audio samples. Similar to the Sound-VECaps dataset, we provide both full-feature captions (AudioCaps-Enhanced) and captions that exclude visual-only contents (AudioCaps-Enhanced) for various evaluation purposes. We also conduct ablation studies to evaluate the effectiveness and the potential of the AudioCaps-Enhanced dataset, which is discussed in the following sections.
4.2 Results
We train all models for 100K steps using the same hyperparameters provided in the paper [14] on Sound-VECaps. AudioLDM-Proposed maintains the same size as AudioLDM [14], while AudioLDM-Proposed-L is a larger version with increased hidden sizes. As shown in Table 2, the proposed AudioLDM achieves SoTA performances on the AudioCaps testing sets. Moreover, the larger model (AudioLDM-Proposed-L), trained on Sound-VECaps, outperforms baseline models by a large margin. As mentioned earlier, current audio generation models struggle with complex and extended prompts, resulting in notable performance degradation on AudioCaps-Enhanced (e.g., the FAD score increases to on AudioLDM2-Large). By using the proposed training dataset, the AudioLDM-Proposed models successfully overcome this limitation, achieving a FAD score of and a MOS score of with the larger version (AudioLDM-Proposed-L).
4.3 Ablation Studies
| Training Dataset | Testing Dataset | KL | IS | FAD |
| Sound-VECaps | AudioCaps | |||
| Sound-VECaps | AudioCaps-E | |||
| Sound-VECaps | AudioCaps-E | |||
| Sound-VECaps | AudioCaps | 8.77 | ||
| Sound-VECaps | AudioCaps-E | 1.17 | ||
| Sound-VECaps | AudioCaps-E | 0.96 |
| Model | Training Set | AudioCaps | AudioCaps-Enhanced | ||||||||||
| Text-to-Audio | Audio-to-Text | Text-to-Audio | Audio-to-Text | ||||||||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| CLAPl [1] | AC+CL+LA | ||||||||||||
| CLAPm [24] | 4.6M-Audio | ||||||||||||
| WavCaps [13] | WavCaps+AC+CL | ||||||||||||
| Auto-ACD [9] | Auto-ACD | 75.3 | 87.4 | ||||||||||
| Sound-VECaps | Sound-VECaps-Audio | 41.2 | |||||||||||
| Sound-VECaps | Sound-VECaps-Full | 54.0 | 85.5 | 93.2 | 53.1 | 85.7 | 91.3 | 64.3 | 90.2 | 96.4 | |||
Effectiveness of Visual-feature. To evaluate the effectiveness of the visual information, we compare the performance of different AudioLDM-Proposed-L systems trained and evaluated on various datasets that include and exclude visual-only content. Notably, all three versions of the testing dataset share the same audio clip for evaluation, hence providing reliability assurance for parallel comparison. As shown in Table 3, the system utilizing Sound-VECaps demonstrates enhanced performance across most evaluation metrics. For the evaluation, using AudioCaps as the prompt presents a higher quality (IS score), while prompts containing visual content exhibit minor degradation. However, in terms of content relevance and audio fidelity, the enriched captions lead to significant improvements, with prompts excluding visual-only content showing SoTA performance. Through these experiments, we have summarized three key findings: 1). Captions with visual features can guide the model toward a better understanding of auditory information, enabling the model to identify features across different modalities; 2). The simplicity of the prompts in current benchmarks limits the presentation of detailed features. The proposed new benchmark with enhanced captions enriches the information and offers greater potential for improving overall performance; 3). Although training with both visual and audio features can provide better results, The additional visual information may increase data complexity during inference. Therefore, system training on prompts without visual-only features (AudioCaps-Enhanced) achieve the best result (FAD score of ) .
Effectiveness on Audio-caption Retrieval.
In addition to our experiments on audio generation, we have assessed the effectiveness of the proposed dataset on audio-language retrieval tasks. Specifically, we employed the WavCaps [13] framework to train and evaluate CLAP-based models in audio retrieval tasks. As illustrated by Table 4, the evaluation using AudioCaps testing set demonstrated that CLAP-based models trained on the Sound-VECaps dataset matched the performance of the baseline models. However, when testing with the enriched and enhanced captions (AudioCaps-Enhanced), the experiment shows a notable performance decline in current SoTA systems, highlighting the challenges posed by longer and more detailed textual features. Conversely, systems trained with enriched captions (Auto-ACD and Sound-VECaps), present improvements in retrieval capabilities, while our proposed dataset achieves the best performance. The results underscore the enhancement of caption through visual information, as well as the accuracy and robustness of the system with our dataset. Additionally, the CLAP model trained with Sound-VECaps exhibited better performance, particularly on feature-enhanced prompts, indicating that the systems can further improve the overall performance by effectively understanding and incorporating these visual-only features.
Effectiveness on Temporal Feature. Despite the enhancement of audio features, another distinguishing aspect of Sound-VECaps is the temporal feature of the captions. Since visual guidance is provided by frame, external temporal information is also included. In detail, we applied the T-Classify method from T-CLAP [25] for temporal feature retrieval. The results in Table 5 demonstrate a stronger capability to identify temporal information in our systems, illustrating the enhanced capability of temporal features contributed by our dataset. Unlike the audio generation and audio retrieval systems, the system developed with audio-only prompts presents better performance than training with full features. This indicates that extensive visual features might influence the model’s understanding of temporal information.
Limitation. We also attempt to use the proposed dataset for several other audio-related tasks. However, due to the rich detailed content in our captions, particularly regarding visual information, the model did not perform well on tasks that are purely audio-targeted, such as audio captioning and zero-shot tasks. These results demonstrate that Sound-VECaps is not a dataset that can be broadly applied to audio-language downstream tasks. It is mainly effective in enhancing the performance of tasks that require processing and distinguishing detailed content.
5 Conclusion
This paper presents a large audio-caption dataset to address the challenge of prompt following in audio generation systems. Specifically, we propose Sound-VECaps, a large-scale dataset comprising over 1.66 million audio clips with captions enriched by visual feature guidance. To evaluate its effectiveness, we developed new benchmarks using AudioCaps testing sets with improved and enriched captions.
Our AudioLDM model with Sound-VECaps achieves state-of-the-art performance and outperforms baseline models. In addition, our system further improves by a large margin when generating complex and enhanced prompts, reaching a FAD score of . The proposed datasets demonstrate substantial improvements in audio retrieval and temporal feature identification. Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simpler prompts and emphasize the value and potential of the proposed benchmark in advancing audio-language model performance. For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed both audio-only versions and captions with full features for different purposes and tasks. Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.
Appendix A Appendix
A.1 LLMs Prompts
We provide the prompts used as the input for the Llama3 model to generate our proposed captions. As shown in the Figure 2, the prompt is a combination of three different features. In the system section, both the caption from Enclap and the audio label are provided, while the frame captions are presented as the user input. Two different contents are also provided for both the full-featured caption (section in green boundaries) and the caption that filtered all the visual-only contents (section in red boundaries). For the AudioCaps-Enhanced dataset, we apply the same prompting pipeline, while changing the caption of enclap into the actual caption provided by the AudioCaps testing set. Nevertheless, all the captions for AudioCaps-Enhanced are generated under human-involved supervision, to ensure the correctness and relevance of the prompts.
A.2 Caption Demos
We present the comparison of the captions from Sound-VECaps and other baseline datasets. As shown in Table 6, for each audio sample, we compare the caption from the AudioSet label, Wavcaps, Enclap, Auto-ACD and two versions of the proposed Sound-VECaps.
Nevertheless, we also present a sample of the proposed AudioCaps-Enhanced testing dataset in Table 7
| Num | Dataset | Caption |
| No.1 | AudioSet | Honk , Speech |
| WavCaps | Crinkling, wind, laughter, ducks, and people speaking are heard. | |
| Enclap | Wind blows, ducks quack and people speak. | |
| Auto-ACD | The wind blows as ducks quack and a man speaks. | |
| VECaps | A goose quacks and honks, while the wind blows, and the person speaks, followed by the sound of bread being offered to the goose, amidst the scattered leaves and grass. | |
| VECaps | As the person stands near the car, a goose quacks and honks, while the wind blows, and the person speaks, followed by the sound of bread being offered to the goose, and the goose’s orange beak and feet can be seen amidst the scattered leaves and grass. | |
| No.2 | AudioSet | Dial tone |
| WavCaps | A dial tone is heard. | |
| Enclap | A telephone rings | |
| Auto-ACD | A dial tone rings with a probability of 0.66, indicating a telephone call in an indoor setting. | |
| VECaps | A telephone rings in the background, followed by a dial tone, while a man is holding a child in his arms, as a news article plays in the background. | |
| VECaps | A telephone rings in the background, followed by a dial tone, while a man is holding a child in his arms in front of a destroyed building, as a news article about US urging Israel to protect civilians and increase aid to Gaza plays in the background. | |
| No.3 | AudioSet | Music, instrument, string |
| WavCaps | Music is playing. | |
| Enclap | A man speaks over a loudspeaker as music plays in the distance | |
| Auto-ACD | The sitar player strums melodious music on stage, accompanied by instruments in an orchestra pit. | |
| VECaps | A man plays a sitar, accompanied by the sound of a plucked string instrument, followed by the soft hum of a bowed string instrument, in a dimly lit room, with music playing in the distance. | |
| VECaps | A man plays the sitar, a traditional Indian stringed instrument, in a dimly lit room with a projection screen in the background, while music plays in the distance, accompanied by the sound of a plucked string instrument, followed by the soft hum of a bowed string instrument. |
| Dataset | Caption |
| AudioCaps | A man talking as water splashes. |
| AudioCaps-E | Waves crashing onto a calm shore, followed by a man speaking amid a gathering of people, some with cameras, by a coastal backdrop. |
| AudioCaps-E | Waves gently lap against the shore under an overcast sky, as a man in a grey shirt and glasses addresses a gathering. Surrounding him, a few individuals, possibly security or journalists, hold cameras and microphones, suggesting a public event near a tropical waterfront. |

### v3

@@
-IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTION
+Sound-VECaps: IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTIONS
 Abstract
@@
 However, existing models struggle with complex and detailed prompts, leading to potential performance degradation.
-We hypothesize that this problem stems from the low quality and relatively small quantity of training data.
-In this work, we aim to create a large-scale audio dataset with rich captions for improving audio generation models.
-We develop an automated pipeline to generate detailed captions for audio-visual datasets by transforming predicted visual captions, audio captions, and tagging labels into comprehensive descriptions using a Large Language Model (LLM).
-We introduce Sound-VECaps, a dataset comprising M high-quality audio-caption pairs with enriched details including audio event orders, occurred places and environment information.
-We demonstrate that training with Sound-VECaps significantly enhances the capability of text-to-audio generation models to comprehend and generate audio from complex input prompts, improving overall system performance.
-Furthermore, we conduct ablation studies of Sound-VECaps across several audio-language tasks, suggesting its potential in advancing audio-text representation learning.
+We hypothesize that this problem stems from the simplicity and scarcity of the training data.
+This work aims to create a large-scale audio dataset with rich captions for improving audio generation models.
+We first develop an automated pipeline to generate detailed captions by transforming predicted visual captions, audio captions, and tagging labels into comprehensive descriptions using a Large Language Model (LLM).
+The resulting dataset, Sound-VECaps, comprises M high-quality audio-caption pairs with enriched details including audio event orders, occurred places and environment information.
+We then demonstrate that training the text-to-audio generation models with Sound-VECaps significantly improves the performance on complex prompts.
+Furthermore, we conduct ablation studies of the models on several downstream audio-language tasks, showing the potential of Sound-VECaps in advancing audio-text representation learning.
 Our dataset and models are available at https://yyua8222.github.io/Sound-VECaps-demo/.
@@
 1 Introduction
-Recent advances in generative modelling have prompted text-to-audio generation models to achieve performance improvements.
-In particular, the development of fundamental language model [1, 2] and sampling-based diffusion model [3, 4] have driven the creation of powerful and robust systems [5, 6, 7] that can generate correlated and high-fidelity audio clips.
-Despite achievements in generating audio with simple and straightforward captions, current models struggle with more complex prompts that present detailed information, which is commonly referred to the challenge as “prompt following ” [8].
-We hypothesize that the fundamental reason for this limitation is the simplicity of current training datasets.
-As highlighted in Table 1, existing audio-caption datasets often fall short in both quality and detailed information.
-The captioning procedures of these datasets mainly rely solely on audio, offering simple captions that average fewer than words per sentence.
-This audio-only approach restricts the ability to provide more useful and enriched content.
-Therefore, the lack of information beyond basic prompts or sound tags restricts the system’s ability to improve its performance on complex and detailed prompts.
-In addition, the simplicity of the caption often results in situations where a single prompt corresponds to multiple audio files, causing instability in the quality and relevance of the generated outputs.
-In contrast, visual features have been proved to convey more information.
-Although the latest large-scale dataset, Auto-ACD [9], applied video features through automatic captioning pipelines to enhance the caption (18 words per caption), it only accounted for a single frame feature, and the prompt has been designed to avoid the visual information.
+Generative models have recently achieved substantial success for text-to-audio generation.
+In particular, the development of language models [1, 2] and diffusion models [3, 4] have enabled the creation of powerful systems [5, 6] on generating high-fidelity audio clips.
+Despite their success in generating audio based on simple captions, current models struggle with complex prompts containing detailed information, which is commonly referred to the challenge as “prompt following” [7].
+A potential reason for this limitation is that existing audio-caption datasets, shown in Table 1, often lack in both quantity and quality (detailed information) of the captions.
+In most of these datasets, each audio is matched with simple and short captions, typically, fewer than words.
+As a result, the captions in these datasets may not contain fine-grained information that could be useful for highly controllable audio generation.
+In addition, the simplicity of the caption often results in situations where the same caption corresponds to multiple audio files (e.g., there are K audio clips match with the caption “ Music is playing” in WavCaps [8]), causing the system to avoid learning specific audio feature and lead to more instability in the generated outputs.
+A possible way to address this issue is to incorporate additional information, such as visual features, which have been shown to provide more detailed insights.
+One of the previous attempts is the Auto-ACD [9], where video features are used to improve the description of the event-occurring scene.
+However, Auto-ACD only takes the visual feature of the middle frame, and the caption has been designed to ignore the visual contents, losing more detailed information.
 | Dataset | Number | Avg. Len | Loc. Inf | Env. Inf |
@@
 | Sound-VECaps | M | M | M |
-In this paper, we aim to leverage external visual information to enhance the captions of the training dataset, guiding the system to follow the prompts.
-The integration of visual and audio features in the caption is designed to train the system to mimic human-like behaviour by constructing complete scenarios within a scene, thereby enhancing the alignment between the prompt and the sound.
+In this paper, we aim to leverage external visual guidance to enhance the audio captions.
+With the improved captions, we can provide better alignment between the prompt and the sound, thereby improving text-to-audio generation systems.
 Specifically, we propose a new pipeline to construct a large-scale audio-language dataset with vision-enhanced captions.
@@
 By incorporating additional visual information, our method ensures the accuracy of audio details while enhancing the captions with comprehensive content, including temporal, spatial, and contextual elements related to the environment.
-Building on the existing largest audio-visual dataset, AudioSet [10], we introduce a new dataset comprising over million audio-caption pairs, named Sound-VECaps.
-This dataset leverages the enhanced capabilities of vision-language integration to produce more detailed and contextually rich audio captions.
+Building on AudioSet [10], we introduce Sound-VECaps, a large-scale dataset comprising over million audio-caption pairs.
 Using Sound-VECaps as the training dataset, our experiments with the audio generation model, AudioLDM [13], show substantial improvements over baseline models.
-In addition, we propose a new benchmark for text-to-audio generation by constructing an enhanced AudioCaps [12] testing set with human-involved supervision (same audio with better prompts).
-Specifically, the AudioLDM-L model trained on Sound-VECaps achieves a Frechet Audio Distance (FAD) score of on the AudioCaps dataset.
-It further improves to a score of on the AudioCaps-Enhanced testing set, significantly outperforming current SoTA models.
-Moreover, we conduct experiments on Sound-VECaps across various audio-language tasks, demonstrating that systems trained on Sound-VECaps also achieve SoTA performance in several specific audio-domain tasks, such as audio retrieval.
-Nevertheless, we investigate the effectiveness of the visual-only content within the caption and the influence of these features during inference.
+To evaluate the performance on complex and extended prompts, we propose a new benchmark for text-to-audio generation by constructing an enhanced AudioCaps [12] testing set (same audio with better captions) named AudioCaps-Enhanced.
+Specifically, the AudioLDM-Large trained on Sound-VECaps achieves a Frechet Audio Distance (FAD) score of on the AudioCaps.
+It further improves to a score of on AudioCaps-Enhanced, significantly outperforming current SoTA models.
+Moreover, we conduct experiments on Sound-VECaps across various audio-language tasks, demonstrating that systems trained on Sound-VECaps achieve SoTA performance in specific audio-domain tasks, such as audio retrieval.
+We also investigate the effectiveness of the visual-only content within the caption and the impact of these features during inference.
 An external version of Sound-VECaps that excludes all the visual-only information (Sound-VECaps) is also provided for different research purposes.
-The remainder of this paper is organized as follows.
-Section 2 introduces the generation pipeline of the proposed dataset, followed by the introduction of the audio generation system which is trained on our datasets in Section 3.
-Section 4 presents the experimental results which evaluate the re-captioned dataset on audio generation as well as several audio-language tasks.
-Conclusions are given in Section 5.
-2 Dataset
+2 Audio Dataset
 As the generation pipeline displayed in Figure 1, the prompt that LLMs used to construct the caption consists of three different information, visual captions from the video, audio captions from the waveform, and the label taggings provided by the original dataset.
 2.1 Visual information Captioning
-We assume that video features can provide more guidance on the events, such as the position, ordering, acoustic environment and some other details of the audio events.
-Different from other captioning approaches [9] that involve a single frame of the whole video, we utilize the complete video information to enable the model to gain a more detailed understanding of each occurring event.
-However, SoTA audio captioning systems [15, 16] follow a similar strategy to understand the whole video clip by averaging the visual feature of the frames into a single feature dimension, which loses the temporal information.
-Hence, we apply image captioning models to capture the visual information for each frame to keep the temporal information.
-For instance, we follow the visual-caption generation pipeline of Stable Diffusion 3 [17] and apply the CogVLM [18] as the visual feature extraction model.
-To improve the efficiency, we clip each video by second and generate the caption of each frame through CogVLM, e.g., captions for eleven frames for a 10-second audio.
+Video features have been proven to provide more guidance on the events [14], such as the position, ordering, and acoustic environment information.
+Different from previous captioning approaches [9] that involve a single frame of the whole video, our video captioning strategy utilizes the complete video information to secure a more detailed description of the occurring audio events.
+However, SoTA audio captioning systems [15, 14] mainly pool the visual feature of each frame into a single feature dimension, losing the temporal information (order of the events).
+Hence, we apply image captioning models to capture the visual information for each frame to reveal the temporal information.
+We follow the visual-caption generation pipeline of the SoTA image generation model, Stable Diffusion 3 [16] and apply the CogVLM [17] for the visual feature extraction.
+To improve the efficiency, the system only captions the frame of each video by second e.g., captions for a -second audio.
 2.2 Audio Information Captioning
-Unlike other modalities, the audio feature is sometimes not highly correlated to visual information, such as background sound sources and invisible sound sources.
-Two constraints are provided as the audio feature for re-captioning to guide the LLM to generate the captions related to the actual auditory information.
+We found that visual information sometimes may not reflect the actual audio features, such as background sound sources and invisible sound sources.
+Hence, two constraints are provided as the audio information to guide the LLM to understand actual auditory information.
 One is the label feature provided by the original AudioSet dataset, another is a simple audio caption generated by the audio captioning model.
-In our system, we applied the SoTA captioning model, EnCLAP [19], to generate concise and brief captions.
+In our system, we applied the SoTA captioning model, EnCLAP [18], to generate the concise and brief captions of each audio clip.
 2.3 Caption Generation
 2.4 Dataset Processing
-Due to the issues of some videos being too old or concerns related to privacy, we collected a total of million videos from the AudioSet dataset.
+Due to the issues of some videos being too old (e.g., not able to be reached and loaded) or concerns related to privacy, we collected a total of M videos from the AudioSet dataset.
 In addition, around k video clips are skipped during the captioning pipeline due to the sensitive policy of the LLMs (e.g., violence).
-Furthermore, we found that some video clips present static visual information and the auditory features are usually background sounds.
-Hence, a filtering strategy is developed to avoid the captions for static video scenes (more than % same frames in each video) and improve the correctness of the dataset.
-Overall, we construct the proposed Sound-VECaps consisting of M audio-caption pairs.
-The Sound-VECaps provides two different versions of captions for different purposes, specifically, Sound-VECaps containing only audible contents or environmental-descriptive information, and Sound-VECaps with full detailed information including visual features, e.g., texts, names, and colours.
+Furthermore, we found that some video clips present static visual information and the auditory features are usually background sounds (irrelevant to visual information).
+To secure the correctness of the visual guidance and improve the quality of the dataset, a filtering strategy is developed to avoid the captions for static video scenes with more than % same frames.
+Overall, we construct the proposed Sound-VECaps datasets consisting of M audio-caption pairs.
+The Sound-VECaps provides two different versions of captions for various purposes, specifically, Sound-VECaps removes visual-only information and contains only audible contents or environmental-descriptive information, while Sound-VECaps describes full detailed information including visual features, e.g., texts, names, and colours.
 | Model | Training Dataset | AudioCaps | AudioCaps-Enhanced | Best Result | |||||
@@
 | AudioLDM2-Large [5] | AC+AS+6 others | 1.22 | |||||||
-| AudioLDM-Proposed | Sound-VECaps | ||||||||
-| AudioLDM-Proposed-L | Sound-VECaps | 1.49 | 1.17 | 7.96 | 1.06 | 43.59 | 4.05 | ||
+| AudioLDM-T5 | Sound-VECaps | ||||||||
+| AudioLDM-T5-L | Sound-VECaps | 1.49 | 1.17 | 7.96 | 1.06 | 43.59 | 4.05 | ||
 3 System Architecture
-To evaluate the impact of the enhanced prompt, we conduct experiments on basic AudioLDM [14] models.
-AudioLDM uses an encoder-decoder architecture to compress the waveform and a diffusion-based generative module to produce the audio features.
-For instance, the system can be divided into four sections: a CLAP encoder for condition embedding, a latent diffusion-based model to generate audio features within the latent space, a variational autoencoder (VAE) decoder to reconstruct the information into a mel spectrogram, and a generative adversarial network (GAN) vocoder [22] to produce the waveform as the final output.
-Instead of using CLAP [1] for audio and text feature embedding during the training and inference stages respectively, our experiment only takes caption as the condition throughout the whole stage.
-Thus the CLAP encoder is replaced with T5 [2] text encoder to understand the textual information better.
-Following the idea of taking the T5 embedding in various lengths with masking [23], a cross-attention module is applied after each convolution layer of the UNet structure to undertake the condition.
-The rest modules exactly follow the design of AudioLDM and our system takes the pre-trained VAE decoder and Hifi-GAN vocoder for audio feature reconstruction.
+To evaluate the impact of the proposed dataset, we conduct experiments on AudioLDM [13] models.
+AudioLDM uses a diffusion-based generative module to produce the audio features and an encoder-decoder architecture to generate the waveform.
+For instance, the system is divided into four sections: a CLAP encoder for condition embedding, a latent diffusion-based model to generate audio features within the latent space, a variational autoencoder (VAE) decoder to reconstruct the information into a mel spectrogram, and a generative adversarial network (HiFi-GAN) vocoder [21] to produce the waveform as the final output.
+Instead of using CLAP [1] for audio and text feature embedding during the training and inference stages respectively, our experiment only takes the caption as the condition throughout the whole stage.
+Thus, the CLAP encoder is replaced with a T5 [2] text encoder to embed the textual information.
+In addition, the across-attention module [22] is applied to undertake the T5 embedding instead of the previous film conditioning module [13], and we name the system as AudioLDM-T5.
+For the remaining modules, we follow the same design of AudioLDM and our system takes the pre-trained VAE decoder and Hifi-GAN vocoder for audio feature reconstruction.
 4 Experiments
 4.1 Evaluation Dataset
-Despite the evaluation procedure on AudioCaps testing set through previous generative models, we introduce a novel benchmark for audio generation with enriched and enhanced captions.
+We first follow previous baseline models and evaluate the performance on the AudioCaps testing set.
+However, AudioCaps only includes simple and audio-only textual information, to better evaluate the system on complex and extended prompts, we introduce a novel benchmark for audio generation with enriched and enhanced captions (same audio with better captions).
 Utilizing the same audio samples, we apply the proposed re-captioning pipeline to generate improved captions for the proposed testing dataset.
-Specifically, human supervision is integrated into the captioning process to ensure the accuracy and semantic relevance of each prompt.
-Following the structure of the original AudioCaps testing set, the proposed AudioCaps-Enhanced testing dataset includes five different captions for each audio clip, totalling captions for audio samples.
+Specifically, human supervision is applied during the captioning process to ensure the accuracy of each caption.
+Same as AudioCaps testing set, the proposed AudioCaps-Enhanced testing dataset includes five different captions for each audio clip, totalling captions for audio samples.
 Similar to the Sound-VECaps dataset, we provide both full-feature captions (AudioCaps-Enhanced) and captions that exclude visual-only contents (AudioCaps-Enhanced) for various evaluation purposes.
@@
 4.2 Results
-We train all models for 100K steps using the same hyperparameters provided in the paper [14] on Sound-VECaps.
-AudioLDM-Proposed maintains the same size as AudioLDM [14], while AudioLDM-Proposed-L is a larger version with increased hidden sizes.
-As shown in Table 2, the proposed AudioLDM achieves SoTA performances on the AudioCaps testing sets.
-Moreover, the larger model (AudioLDM-Proposed-L), trained on Sound-VECaps, outperforms baseline models by a large margin.
-As mentioned earlier, current audio generation models struggle with complex and extended prompts, resulting in notable performance degradation on AudioCaps-Enhanced (e.g., the FAD score increases to on AudioLDM2-Large).
-By using the proposed training dataset, the AudioLDM-Proposed models successfully overcome this limitation, achieving a FAD score of and a MOS score of with the larger version (AudioLDM-Proposed-L).
-4.3 Ablation Studies
+Effectiveness on Audio Generation.
+All the models are trained for 100K steps on Sound-VECaps using the same hyperparameters of AudioLDM.
+Specifically, AudioLDM-T5 maintains the same size as AudioLDM [13], while AudioLDM-T5-L is a larger version system with increased hidden sizes.
+As shown in Table 2, the AudioLDM-T5 achieves SoTA performance on the AudioCaps testing sets.
+Moreover, the larger model (AudioLDM-T5-L), trained on Sound-VECaps, outperforms baseline models by a large margin.
+In addition, current audio generation models struggle with complex and extended prompts, resulting in notable performance degradation on AudioCaps-Enhanced (e.g., the FAD score increases from to on AudioLDM2-Large).
+By applying Sound-VECaps as a training dataset, the AudioLDM-T5 models successfully overcome this limitation, achieving a FAD score of and a MOS score of with the larger version (AudioLDM-T5-L).
 | Training Dataset | Testing Dataset | KL | IS | FAD |
@@
 Effectiveness of Visual-feature.
-To evaluate the effectiveness of the visual information, we compare the performance of different AudioLDM-Proposed-L systems trained and evaluated on various datasets that include and exclude visual-only content.
-Notably, all three versions of the testing dataset share the same audio clip for evaluation, hence providing reliability assurance for parallel comparison.
-As shown in Table 3, the system utilizing Sound-VECaps demonstrates enhanced performance across most evaluation metrics.
-For the evaluation, using AudioCaps as the prompt presents a higher quality (IS score), while prompts containing visual content exhibit minor degradation.
-However, in terms of content relevance and audio fidelity, the enriched captions lead to significant improvements, with prompts excluding visual-only content showing SoTA performance.
+To evaluate the effectiveness of the visual information in the captions, we compare the performance of different AudioLDM-T5-L systems trained and evaluated on various datasets that include and exclude visual-only content.
+Notably, all three versions of the testing dataset share the same group of audio clips (same target audio samples while using different prompts for generation), providing reliability assurance for the comparison.
+As shown in Table 3, systems utilizing Sound-VECaps as the training dataset demonstrates enhanced performance across all three evaluation metrics.
+For the evaluation, using AudioCaps as the caption presents a higher quality (IS score of ), while the audio outputs generated through the captions with visual content (AudioCaps-Enhanced) show minor degradation.
+However, audio samples generated through enriched prompts lead to significant improvements in the fidelity of generated audio, with the prompts excluding visual-only content (AudioCaps-Enhanced) showing SoTA performance.
 Through these experiments, we have summarized three key findings: 1).
-Captions with visual features can guide the model toward a better understanding of auditory information, enabling the model to identify features across different modalities; 2).
-The simplicity of the prompts in current benchmarks limits the presentation of detailed features.
-The proposed new benchmark with enhanced captions enriches the information and offers greater potential for improving overall performance; 3).
-Although training with both visual and audio features can provide better results, The additional visual information may increase data complexity during inference.
-Therefore, system training on prompts without visual-only features (AudioCaps-Enhanced) achieve the best result (FAD score of ) .
-Effectiveness on Audio-caption Retrieval.
-In addition to our experiments on audio generation, we have assessed the effectiveness of the proposed dataset on audio-language retrieval tasks.
-Specifically, we employed the WavCaps [13] framework to train and evaluate CLAP-based models in audio retrieval tasks.
-As illustrated by Table 4, the evaluation using AudioCaps testing set demonstrated that CLAP-based models trained on the Sound-VECaps dataset matched the performance of the baseline models.
-However, when testing with the enriched and enhanced captions (AudioCaps-Enhanced), the experiment shows a notable performance decline in current SoTA systems, highlighting the challenges posed by longer and more detailed textual features.
-Conversely, systems trained with enriched captions (Auto-ACD and Sound-VECaps), present improvements in retrieval capabilities, while our proposed dataset achieves the best performance.
-The results underscore the enhancement of caption through visual information, as well as the accuracy and robustness of the system with our dataset.
-Additionally, the CLAP model trained with Sound-VECaps exhibited better performance, particularly on feature-enhanced prompts, indicating that the systems can further improve the overall performance by effectively understanding and incorporating these visual-only features.
-Effectiveness on Temporal Feature.
-Despite the enhancement of audio features, another distinguishing aspect of Sound-VECaps is the temporal feature of the captions.
-Since visual guidance is provided by frame, external temporal information is also included.
-In detail, we applied the T-Classify method from T-CLAP [25] for temporal feature retrieval.
-The results in Table 5 demonstrate a stronger capability to identify temporal information in our systems, illustrating the enhanced capability of temporal features contributed by our dataset.
+Training on captions with visual features can improve the capability of the system to handle auditory information and identify features across different modalities, leading to significant improvement in the overall performance; 2).
+The simplicity of the prompts in current evaluation benchmarks (e.g.
+AudioCaps) limits the presentation of detailed audio features.
+The proposed benchmark testing on AudioCaps-Enhanced enriches the information with more controllable features and offers greater potential for enhancing the output quality; 3).
+Although training with external visual features (Sound-VECaps) provides better results, the additional visual information may increase data complexity during inference.
+Therefore, the results generated on prompts without visual-only features (AudioCaps-Enhanced) further achieve the best result (a FAD score of ).
+4.3 Studies on Other Audio-tasks
+Audio-caption Retrieval.
+In addition to our experiments on audio generation, we assessed the effectiveness of Sound-VECaps on audio-language retrieval tasks.
+Specifically, we employed the WavCaps [8] framework (RoBERTa text encoder and HTSAT audio encoder) to train and evaluate CLAP-based models in audio retrieval tasks.
+As illustrated by Table 4, the evaluation using the AudioCaps testing set demonstrated that CLAP-based models trained on the Sound-VECaps dataset matched the performance of the baseline models.
+However, when testing with the enriched and enhanced captions (AudioCaps-Enhanced), the experiment shows a notable performance decline in current SoTA systems, highlighting the challenges posed by longer and more detailed textual information.
+Conversely, systems trained with enriched captions (Auto-ACD and Sound-VECaps) present improvements in retrieval capabilities, while the system on Sound-VECaps achieves the best performance.
+The results show the enhancement of caption through visual information, as well as the accuracy and robustness of the system on Sound-VECaps.
+Additionally, the CLAP model trained with Sound-VECaps exhibited better performance, particularly on AudioCaps-Enhanced, indicating that the systems can further improve the overall performance by effectively understanding and incorporating these visual-only features.
+Temporal Feature Retrieval.
+Another distinguishing aspect of Sound-VECaps is the temporal feature of the captions.
+Since visual guidance is provided by frame, external temporal information is also included in the proposed captions.
+For the evaluation of temporal feature retrieval, we applied the T-Classify method from T-CLAP [24].
+The results in Table 5 demonstrate a stronger capability to identify temporal information in the system on Sound-VECaps, illustrating the enhanced capability of temporal features contributed by the dataset.
 Unlike the audio generation and audio retrieval systems, the system developed with audio-only prompts presents better performance than training with full features.
@@
 We also attempt to use the proposed dataset for several other audio-related tasks.
-However, due to the rich detailed content in our captions, particularly regarding visual information, the model did not perform well on tasks that are purely audio-targeted, such as audio captioning and zero-shot tasks.
+However, due to the rich detailed content in our captions, particularly regarding visual information, the model did not perform well on tasks that are purely audio-targeted content, such as audio captioning and zero-shot tasks.
 These results demonstrate that Sound-VECaps is not a dataset that can be broadly applied to audio-language downstream tasks.
-It is mainly effective in enhancing the performance of tasks that require processing and distinguishing detailed content.
+It is mainly effective in enhancing the performance of tasks that require processing and distinguishing detailed content, such as generation and retrieval.
 5 Conclusion
-This paper presents a large audio-caption dataset to address the challenge of prompt following in audio generation systems.
-Specifically, we propose Sound-VECaps, a large-scale dataset comprising over 1.66 million audio clips with captions enriched by visual feature guidance.
-To evaluate its effectiveness, we developed new benchmarks using AudioCaps testing sets with improved and enriched captions.
-Our AudioLDM model with Sound-VECaps achieves state-of-the-art performance and outperforms baseline models.
-In addition, our system further improves by a large margin when generating complex and enhanced prompts, reaching a FAD score of .
-The proposed datasets demonstrate substantial improvements in audio retrieval and temporal feature identification.
-Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simpler prompts and emphasize the value and potential of the proposed benchmark in advancing audio-language model performance.
-For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed both audio-only versions and captions with full features for different purposes and tasks.
+This paper presents Sound-VECaps, a large-scale dataset comprising over M audio clips with captions enriched by visual feature guidance, to address the challenge of prompt following in audio generation systems.
+In addition, a new benchmark using improved and enriched captions is proposed to evaluate audio-language systems on complex and extended prompts.
+Experiments show that AudioLDM models trained on Sound-VECaps achieve SoTA performance and outperform baseline models.
+In addition, our systems further improve by a large margin when taking more detailed captions as prompts, reaching a FAD score of .
+In addition, Sound-VECaps demonstrate substantial improvements in audio retrieval and temporal feature identification.
+Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simple prompts and emphasize the value and potential of the more detailed prompts in advancing the performance of audio-language models.
+For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed audio-only versions and full versions with captions that include and exclude visual-only content for different purposes and tasks.
 Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.
+6 ACKNOWLEDGMENT
+This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDnace, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”.
 Appendix A Appendix

### v4

@@
-Sound-VECaps: IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTIONS
+Sound-VECaps: Improving Audio Generation with Visually Enhanced Captions
 Abstract
@@
 Furthermore, we conduct ablation studies of the models on several downstream audio-language tasks, showing the potential of Sound-VECaps in advancing audio-text representation learning.
-Our dataset and models are available at https://yyua8222.github.io/Sound-VECaps-demo/.
-Index Terms— Audio generation, audio retrieval, diffusion model, audio-language dataset
-1 Introduction
+Details of our dataset and demos are available here.
+Index Terms:
+Audio generation, audio retrieval, diffusion model, audio-language datasetI Introduction
 Generative models have recently achieved substantial success for text-to-audio generation.
 In particular, the development of language models [1, 2] and diffusion models [3, 4] have enabled the creation of powerful systems [5, 6] on generating high-fidelity audio clips.
-Despite their success in generating audio based on simple captions, current models struggle with complex prompts containing detailed information, which is commonly referred to the challenge as “prompt following” [7].
-A potential reason for this limitation is that existing audio-caption datasets, shown in Table 1, often lack in both quantity and quality (detailed information) of the captions.
+Despite their success in generating audio with simple captions, current models struggle with complex prompts containing detailed information, which referred to the challenge as “prompt following” [3].
+A potential reason for this limitation is that existing audio-caption datasets often lack in quantity and quality (detailed information) of the captions.
 In most of these datasets, each audio is matched with simple and short captions, typically, fewer than words.
@@
 One of the previous attempts is the Auto-ACD [8], where video features are used to improve the description of the event-occurring scene.
-However, Auto-ACD only takes the visual feature of the middle frame, and the caption has been designed to ignore the visual contents, losing more detailed information.
-| Dataset | Number | Avg. Len | Loc. Inf | Env. Inf |
-| AudioSet [10] | M | Label | Label | |
-| Clotho [11] | K | K | K | |
-| AudioCaps [12] | K | K | K | |
-| WavCaps [8] | K | K | K | |
-| Auto-ACD [9] | M | M | K | |
-| Sound-VECaps | M | M | M | |
-| Sound-VECaps | M | M | M |
+However, Auto-ACD only takes the visual feature of the middle frame, and the caption has been designed to ignore the visual-only contents, losing more detailed information.
 In this paper, we aim to leverage external visual guidance to enhance the audio captions.
-With the improved captions, we can provide better alignment between the prompt and the sound, thereby improving text-to-audio generation systems.
-Specifically, we propose a new pipeline to construct a large-scale audio-language dataset with vision-enhanced captions.
+With improved captions, we can provide better alignment between the prompt and the sound, thereby improving text-to-audio generation systems.
+Specifically, we propose new pipelines to construct a large-scale audio-language dataset with vision-enhanced captions.
 Our approach first involves collecting external visual information using state-of-the-art (SoTA) image captioning models.
-These visual captions, combined with audio information, are then used to create new, enriched captions through Large Language Models (LLMs).
+These visual captions, combined with simple audio information, are then used to create new, enriched captions through Large Language Models (LLMs).
 By incorporating additional visual information, our method ensures the accuracy of audio details while enhancing the captions with comprehensive content, including temporal, spatial, and contextual elements related to the environment.
-Building on AudioSet [10], we introduce Sound-VECaps, a large-scale dataset comprising over million audio-caption pairs.
+Building on AudioSet [9], we introduce Sound-VECaps, a large-scale dataset comprising over M audio-caption pairs.
 Using Sound-VECaps as the training dataset, our experiments with the audio generation model, AudioLDM [12], show substantial improvements over baseline models.
@@
 We also investigate the effectiveness of the visual-only content within the caption and the impact of these features during inference.
-An external version of Sound-VECaps that excludes all the visual-only information (Sound-VECaps) is also provided for different research purposes.
-2 Audio Dataset
-As the generation pipeline displayed in Figure 1, the prompt that LLMs used to construct the caption consists of three different information, visual captions from the video, audio captions from the waveform, and the label taggings provided by the original dataset.
-2.1 Visual information Captioning
-Video features have been proven to provide more guidance on the events [14], such as the position, ordering, and acoustic environment information.
-Different from previous captioning approaches [9] that involve a single frame of the whole video, our video captioning strategy utilizes the complete video information to secure a more detailed description of the occurring audio events.
-However, SoTA audio captioning systems [15, 14] mainly pool the visual feature of each frame into a single feature dimension, losing the temporal information (order of the events).
-Hence, we apply image captioning models to capture the visual information for each frame to reveal the temporal information.
-We follow the visual-caption generation pipeline of the SoTA image generation model, Stable Diffusion 3 [16] and apply the CogVLM [17] for the visual feature extraction.
+In addition, an external version of Sound-VECaps that excludes all the visual-only information (Sound-VECaps) is also provided for different research purposes.
+II Audio Dataset
+Our Sound-VECaps dataset is built on AudioSet [9] with the processing pipeline shown in Figure 1.
+In particular, LLMs are prompted to generate captions based on three pieces of text information, namely, visual captions from the video, audio captions from the waveform, and the label tags in the original dataset.
+II-A Captions from Video
+One of the novel aspects of the proposed dataset is that we leverage the caption of the corresponding video to provide more detailed information about the audio events.
+Different from previous visual-related approaches [8] that only apply the visual information of middle frames, the proposed strategy utilizes the captions of complete videos to secure more detailed descriptions.
+On the other hand, current SoTA video captioning systems [13, 14] mainly pool all the visual information into an aligned feature dimension, losing the temporal information (order of the events).
+Hence, we capture visual information for multiple frames through image captioning to maintain this temporal information.
+Specifically, we follow the image caption generation of Stable Diffusion 3 [15] and apply the CogVLM [16] captioning system.
 To improve the efficiency, the system only captions the frame of each video by second e.g., captions for a -second audio.
-2.2 Audio Information Captioning
-We found that visual information sometimes may not reflect the actual audio features, such as background sound sources and invisible sound sources.
-Hence, two constraints are provided as the audio information to guide the LLM to understand actual auditory information.
-One is the label feature provided by the original AudioSet dataset, another is a simple audio caption generated by the audio captioning model.
-In our system, we applied the SoTA captioning model, EnCLAP [18], to generate the concise and brief captions of each audio clip.
-2.3 Caption Generation
-2.4 Dataset Processing
-Due to the issues of some videos being too old (e.g., not able to be reached and loaded) or concerns related to privacy, we collected a total of M videos from the AudioSet dataset.
-In addition, around k video clips are skipped during the captioning pipeline due to the sensitive policy of the LLMs (e.g., violence).
-Furthermore, we found that some video clips present static visual information and the auditory features are usually background sounds (irrelevant to visual information).
-To secure the correctness of the visual guidance and improve the quality of the dataset, a filtering strategy is developed to avoid the captions for static video scenes with more than % same frames.
-Overall, we construct the proposed Sound-VECaps datasets consisting of M audio-caption pairs.
-The Sound-VECaps provides two different versions of captions for various purposes, specifically, Sound-VECaps removes visual-only information and contains only audible contents or environmental-descriptive information, while Sound-VECaps describes full detailed information including visual features, e.g., texts, names, and colours.
+II-B Captions from Audio
+We found that captions directly from video sometimes may not reflect the correct events, such as background and invisible sound.
+Hence, two constraints are provided to guide the LLM to understand the auditory information: the label provided by the original AudioSet dataset, and the simple audio caption generated by audio captioning models.
+We apply the SoTA captioning model, EnCLAP [17], to generate concise and brief captions of each audio clip.
+II-C Proposed Caption Generation
+Combining three textual information mentioned above, an LLM is applied to generate the final caption, where we use Llama3-7B [18] to assemble re-caption the comprehensive description of each audio.
+II-D Dataset Processing
+Due to the issues of some videos being too old (not accessible anymore), we collected a total of M videos from the AudioSet.
+In addition, around k video clips are skipped due to the sensitive policy of the LLMs (e.g., violence).
+Furthermore, we found that some video clips present static visual information with complete background sounds, leading the caption focusing on visual events but ignoring the actual audio events.
+To ensure the correctness of the visual guidance and improve the data quality, a filtering strategy is applied to detect and exclude the captions of static video which presents more than % same frames.
+Overall, we obtain the Sound-VECaps datasets containing M audio-caption pairs.
+The Sound-VECaps provides two different versions of captions for various purposes, specifically, Sound-VECaps removes visual-only information and contains only audible contents or environmental-descriptive information, while Sound-VECaps describes full detailed information including visual features, e.g., texts, names, shapes, and colours.
 | Model | Training Dataset | AudioCaps | AudioCaps-Enhanced | Best Result | |||||
@@
 | AudioLDM-T5-L | Sound-VECaps | 1.49 | 1.17 | 7.96 | 1.06 | 43.59 | 4.05 | ||
-3 System Architecture
-To evaluate the impact of the proposed dataset, we conduct experiments on AudioLDM [13] models.
-AudioLDM uses a diffusion-based generative module to produce the audio features and an encoder-decoder architecture to generate the waveform.
-For instance, the system is divided into four sections: a CLAP encoder for condition embedding, a latent diffusion-based model to generate audio features within the latent space, a variational autoencoder (VAE) decoder to reconstruct the information into a mel spectrogram, and a generative adversarial network (HiFi-GAN) vocoder [21] to produce the waveform as the final output.
-Instead of using CLAP [1] for audio and text feature embedding during the training and inference stages respectively, our experiment only takes the caption as the condition throughout the whole stage.
-Thus, the CLAP encoder is replaced with a T5 [2] text encoder to embed the textual information.
-In addition, the across-attention module [22] is applied to undertake the T5 embedding instead of the previous film conditioning module [13], and we name the system as AudioLDM-T5.
-For the remaining modules, we follow the same design of AudioLDM and our system takes the pre-trained VAE decoder and Hifi-GAN vocoder for audio feature reconstruction.
-4 Experiments
-4.1 Evaluation Dataset
-We first follow previous baseline models and evaluate the performance on the AudioCaps testing set.
-However, AudioCaps only includes simple and audio-only textual information, to better evaluate the system on complex and extended prompts, we introduce a novel benchmark for audio generation with enriched and enhanced captions (same audio with better captions).
-Utilizing the same audio samples, we apply the proposed re-captioning pipeline to generate improved captions for the proposed testing dataset.
-Specifically, human supervision is applied during the captioning process to ensure the accuracy of each caption.
-Same as AudioCaps testing set, the proposed AudioCaps-Enhanced testing dataset includes five different captions for each audio clip, totalling captions for audio samples.
-Similar to the Sound-VECaps dataset, we provide both full-feature captions (AudioCaps-Enhanced) and captions that exclude visual-only contents (AudioCaps-Enhanced) for various evaluation purposes.
-We also conduct ablation studies to evaluate the effectiveness and the potential of the AudioCaps-Enhanced dataset, which is discussed in the following sections.
-4.2 Results
+III Audio Generation System
+We conduct experiments on text-to-audio generation using AudioLDM [12] models, a SoTA audio generation model, to evaluate Sound-VECaps.
+For instance, AudioLDM is divided into four sections: a CLAP encoder for condition embedding, a latent diffusion-based model to generate audio features within the latent space, a variational autoencoder (VAE) decoder to reconstruct the information into a mel spectrogram, and a generative adversarial network(HiFi-GAN) vocoder [20] to produce the waveform as the final output.
+Instead of using CLAP [1] for computing the audio and text embedding, our experiments replaced the encoder with a T5 [2] encoder for condition embedding, and an cross-attention module [21] is applied to process the T5 embedding.
+We name the system as AudioLDM-T5.
+For the remaining modules, we follow the same design of AudioLDM and our system takes the pre-trained VAE decoder and Hifi-GAN [20] vocoder for audio feature reconstruction.
+IV Experiments
+IV-A Evaluation Dataset
+We first follow previous baseline models [12, 19] and evaluate the performance of text-to-audio generation on the AudioCaps testing set.
+Since AudioCaps only includes simple and audio-only textual information, we introduce a novel benchmark with enriched and enhanced captions (same audio with better captions).
+We apply the proposed re-captioning pipeline in Section II to generate improved captions for AudioCaps testing audio samples.
+Specifically, human supervision is applied during the captioning process to check the accuracy and relevance of each caption (ensure the quality of LLM outputs).
+The proposed AudioCaps-Enhanced testing dataset includes five different captions for each audio clip, totalling captions for audio samples.
+Similar to the Sound-VECaps dataset, we provide both full-feature captions (AudioCaps-Enhanced) and captions excluding visual-only contents (AudioCaps-Enhanced) for various purposes.
+IV-B Results
 Effectiveness on Audio Generation.
-All the models are trained for 100K steps on Sound-VECaps using the same hyperparameters of AudioLDM.
-Specifically, AudioLDM-T5 maintains the same size as AudioLDM [13], while AudioLDM-T5-L is a larger version system with increased hidden sizes.
-As shown in Table 2, the AudioLDM-T5 achieves SoTA performance on the AudioCaps testing sets.
+Audio generation systems are trained on Sound-VECaps to evaluate their effectiveness, where models are trained using the same hyperparameters of AudioLDM.
+Specifically, AudioLDM-T5 maintains the same size as AudioLDM [12], while AudioLDM-T5-L is a larger system with increased hidden sizes.
+As shown in Table II, AudioLDM-T5 achieves SoTA performance on the AudioCaps testing sets.
 Moreover, the larger model (AudioLDM-T5-L), trained on Sound-VECaps, outperforms baseline models by a large margin.
-In addition, current audio generation models struggle with complex and extended prompts, resulting in notable performance degradation on AudioCaps-Enhanced (e.g., the FAD score increases from to on AudioLDM2-Large).
-By applying Sound-VECaps as a training dataset, the AudioLDM-T5 models successfully overcome this limitation, achieving a FAD score of and a MOS score of with the larger version (AudioLDM-T5-L).
+In addition, current audio generation models struggle with complex and extended prompts, resulting in performance degradation on AudioCaps-Enhanced (e.g., the FAD score increases from to on AudioLDM2-L).
+By applying Sound-VECaps for training, AudioLDM-T5 models successfully overcome this limitation, achieving a FAD score of and a MOS score of on larger AudioLDM-T5-L.
 | Training Dataset | Testing Dataset | KL | IS | FAD |
@@
 | Auto-ACD [8] | Auto-ACD | 75.3 | 87.4 | ||||||||||
-| Sound-VECaps | Sound-VECaps-Audio | 41.2 | |||||||||||
-| Sound-VECaps | Sound-VECaps-Full | 54.0 | 85.5 | 93.2 | 53.1 | 85.7 | 91.3 | 64.3 | 90.2 | 96.4 | |||
-Effectiveness of Visual-feature.
+| CLAP | Sound-VECaps-Audio | 41.2 | |||||||||||
+| CLAP | Sound-VECaps-Full | 54.0 | 85.5 | 93.2 | 53.1 | 85.7 | 91.3 | 64.3 | 90.2 | 96.4 | |||
+Effectiveness of Visual-Only Content.
 To evaluate the effectiveness of the visual information in the captions, we compare the performance of different AudioLDM-T5-L systems trained and evaluated on various datasets that include and exclude visual-only content.
 Notably, all three versions of the testing dataset share the same group of audio clips (same target audio samples while using different prompts for generation), providing reliability assurance for the comparison.
-As shown in Table 3, systems utilizing Sound-VECaps as the training dataset demonstrates enhanced performance across all three evaluation metrics.
-For the evaluation, using AudioCaps as the caption presents a higher quality (IS score of ), while the audio outputs generated through the captions with visual content (AudioCaps-Enhanced) show minor degradation.
+As shown in Table III, systems utilizing Sound-VECaps as the training dataset demonstrates enhanced performance across all three evaluation metrics.
+For the evaluation, using AudioCaps as the prompt presents a higher quality (IS score of ), while the audio outputs generated through the prompts with visual content (AudioCaps-Enhanced) show minor degradation.
 However, audio samples generated through enriched prompts lead to significant improvements in the fidelity of generated audio, with the prompts excluding visual-only content (AudioCaps-Enhanced) showing SoTA performance.
-Through these experiments, we have summarized three key findings: 1).
-Training on captions with visual features can improve the capability of the system to handle auditory information and identify features across different modalities, leading to significant improvement in the overall performance; 2).
-The simplicity of the prompts in current evaluation benchmarks (e.g.
+Through these experiments, we have summarized three key findings: 1) Training on captions with visual features can improve the capability of the system to handle auditory information and identify features across different modalities, leading to significant improvement in the overall performance.
+2) The simplicity of the prompts in current evaluation benchmarks (e.g.
 AudioCaps) limits the presentation of detailed audio features.
-The proposed benchmark testing on AudioCaps-Enhanced enriches the information with more controllable features and offers greater potential for enhancing the output quality; 3).
-Although training with external visual features (Sound-VECaps) provides better results, the additional visual information may increase data complexity during inference.
-Therefore, the results generated on prompts without visual-only features (AudioCaps-Enhanced) further achieve the best result (a FAD score of ).
-4.3 Studies on Other Audio-tasks
-Audio-caption Retrieval.
-In addition to our experiments on audio generation, we assessed the effectiveness of Sound-VECaps on audio-language retrieval tasks.
-Specifically, we employed the WavCaps [8] framework (RoBERTa text encoder and HTSAT audio encoder) to train and evaluate CLAP-based models in audio retrieval tasks.
-As illustrated by Table 4, the evaluation using the AudioCaps testing set demonstrated that CLAP-based models trained on the Sound-VECaps dataset matched the performance of the baseline models.
-However, when testing with the enriched and enhanced captions (AudioCaps-Enhanced), the experiment shows a notable performance decline in current SoTA systems, highlighting the challenges posed by longer and more detailed textual information.
-Conversely, systems trained with enriched captions (Auto-ACD and Sound-VECaps) present improvements in retrieval capabilities, while the system on Sound-VECaps achieves the best performance.
-The results show the enhancement of caption through visual information, as well as the accuracy and robustness of the system on Sound-VECaps.
-Additionally, the CLAP model trained with Sound-VECaps exhibited better performance, particularly on AudioCaps-Enhanced, indicating that the systems can further improve the overall performance by effectively understanding and incorporating these visual-only features.
+The proposed benchmark testing on AudioCaps-Enhanced enriches the information with more controllable features and offers greater potential for enhancing the output quality.
+3) Although training with external visual features (Sound-VECaps) provides better results, the additional visual information may increase data complexity during inference.
+Therefore, the system that uses prompts without visual-only features (AudioCaps-Enhanced) generates the best result ( on FAD).
+IV-C Studies on Other Audio Tasks
+Audio Caption Retrieval.
+In addition to our experiments on audio generation, we evaluated the effectiveness of Sound-VECaps for improving audio-language retrieval systems.
+Specifically, we employed the framework in WavCaps [7], which uses RoBERTa as the text encoder and HTSAT as the audio encoder, to train and evaluate CLAP-based models in audio-text cross modal retrieval tasks.
+As illustrated by Table IV, the evaluation using the AudioCaps testing set demonstrated that the CLAP-based models trained on the Sound-VECaps dataset matched the performance of the baseline models (trained on other datasets).
+However, when testing with the enhanced captions (AudioCaps-Enhanced), the experiment shows a notable performance decline in current SoTA systems, highlighting the challenges posed by longer and more detailed textual information.
+Conversely, the systems trained with enriched captions (Auto-ACD [8] and Sound-VECaps) present improvements in retrieval capabilities, while the system on Sound-VECaps achieves the best performance.
+The results show the enhancement of captions through visual information, as well as the accuracy and robustness of the system on Sound-VECaps.
+Additionally, the CLAP model trained with Sound-VECaps exhibited better performance, particularly on AudioCaps-Enhanced dataset, indicating that the overall performance of the system can be further improved with the visually augmented captions.
 Temporal Feature Retrieval.
-Another distinguishing aspect of Sound-VECaps is the temporal feature of the captions.
-Since visual guidance is provided by frame, external temporal information is also included in the proposed captions.
-For the evaluation of temporal feature retrieval, we applied the T-Classify method from T-CLAP [24].
-The results in Table 5 demonstrate a stronger capability to identify temporal information in the system on Sound-VECaps, illustrating the enhanced capability of temporal features contributed by the dataset.
-Unlike the audio generation and audio retrieval systems, the system developed with audio-only prompts presents better performance than training with full features.
-This indicates that extensive visual features might influence the model’s understanding of temporal information.
+Another aspect of Sound-VECaps is the temporal information.
+Since visual captions are provided by frame, temporal information (events ordering) is also included.
+We applied the T-Classify method from T-CLAP [23] to evaluate the performance of temporal feature retrieval.
+Table V demonstrates stronger capabilities to identify temporal information in systems with Sound-VECaps, illustrating its improvement in temporal features.
+In addition, the system developed without visual-only contents presents better performance, indicating that extensive visual features might influence the model’s understanding of temporal information.
 Limitation.
-We also attempt to use the proposed dataset for several other audio-related tasks.
-However, due to the rich detailed content in our captions, particularly regarding visual information, the model did not perform well on tasks that are purely audio-targeted content, such as audio captioning and zero-shot tasks.
-These results demonstrate that Sound-VECaps is not a dataset that can be broadly applied to audio-language downstream tasks.
-It is mainly effective in enhancing the performance of tasks that require processing and distinguishing detailed content, such as generation and retrieval.
-5 Conclusion
-This paper presents Sound-VECaps, a large-scale dataset comprising over M audio clips with captions enriched by visual feature guidance, to address the challenge of prompt following in audio generation systems.
-In addition, a new benchmark using improved and enriched captions is proposed to evaluate audio-language systems on complex and extended prompts.
-Experiments show that AudioLDM models trained on Sound-VECaps achieve SoTA performance and outperform baseline models.
-In addition, our systems further improve by a large margin when taking more detailed captions as prompts, reaching a FAD score of .
-In addition, Sound-VECaps demonstrate substantial improvements in audio retrieval and temporal feature identification.
-Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simple prompts and emphasize the value and potential of the more detailed prompts in advancing the performance of audio-language models.
-For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed audio-only versions and full versions with captions that include and exclude visual-only content for different purposes and tasks.
-Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.
-6 ACKNOWLEDGMENT
-This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDnace, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”.
+We also attempt to use Sound-VECaps for other audio-related tasks.
+However, due to the rich content in our captions, particularly regarding visual information, the model did not perform well on tasks that are purely audio-targeted content, such as audio captioning and zero-shot tasks.
+These results demonstrate that Sound-VECaps may not be broadly applied to audio-language tasks.
+It is mainly effective in a range of tasks that require processing and distinguishing detailed content, such as generation and retrieval.
+V Conclusion
+We present Sound-VECaps, a large-scale dataset comprising M audio clips with captions augmented by video data, to address the challenge of prompt following in audio generation systems.
+Experiments show that systems trained on Sound-VECaps achieve SoTA performance and outperform baseline models.
+In addition, a new benchmark using improved captions is proposed to evaluate audio-language systems on complex and extended prompts.
+Our systems are further improved by a large margin when taking more detailed captions as prompts, reaching a FAD score of .
+Nevertheless, we demonstrated that using Sound-VECaps can offer substantial improvements in audio-language and temporal feature retrieval.
+The results of the proposed AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks and emphasize the potential of better prompts in advancing the performance of audio-language models.
+Overall, we develop two versions of the proposed datasets with captions that include and exclude visual-only content for different purposes and tasks and hope these datasets will generate more profound impacts on audio-language learning.
+VI ACKNOWLEDGMENT
+This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDance, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”.
+For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
 Appendix A Appendix
-A.1 LLMs Prompts
+A-A LLMs Prompts
 We provide the prompts used as the input for the Llama3 model to generate our proposed captions.
@@
 Nevertheless, all the captions for AudioCaps-Enhanced are generated under human-involved supervision, to ensure the correctness and relevance of the prompts.
-A.2 Caption Demos
+A-B Caption Demos
 We present the comparison of the captions from Sound-VECaps and other baseline datasets.
-As shown in Table 6, for each audio sample, we compare the caption from the AudioSet label, Wavcaps, Enclap, Auto-ACD and two versions of the proposed Sound-VECaps.
-Nevertheless, we also present a sample of the proposed AudioCaps-Enhanced testing dataset in Table 7
+As shown in Table VI, for each audio sample, we compare the caption from the AudioSet label, Wavcaps, Enclap, Auto-ACD and two versions of the proposed Sound-VECaps.
+A sample of the proposed AudioCaps-Enhanced testing dataset is presented in Table VII.
 | Num | Dataset | Caption |